OpenStudy (anonymous):

[(1/x) - (1/y)] / [(1/x) + (1/y)] Help?

OpenStudy (anonymous):

Explain pleasee.?

OpenStudy (anonymous):

Common Denominator: xy y/xy - x/xy (x - y)/xy TOP (x + y)/xy Bottom Flip to multiply (x-y)/(xy) (xy)/(x+y) (y-x)/(x+y)

OpenStudy (amistre64):

I usually go an algebra route and do this: then solve and simplify for N \[\frac{(1/x) - (1/y)}{(1/x) + (1/y)}=N\]

OpenStudy (akshay_budhkar):

hey! my answer?

OpenStudy (anonymous):

Interesting strategy amistre64, I use that to solve for logarithms with odd bases and no calculator.

OpenStudy (anonymous):

When working with complex fractions, I prefer to simplify them first. LCD of the inner fractions is xy \[\frac{(1/x) - (1/y)}{(1/x) + (1/y)}\]\[=\frac{(1/x) - (1/y)}{(1/x) + (1/y)} \cdot \frac{xy}{xy}\]\[=\frac{y - x}{y + x}\]

OpenStudy (anonymous):

So whenever you have a complex fraction you can simplify by multiplying the top/bottom of the outer (larger) fraction by the LCD of all the inner fractions

OpenStudy (amistre64):

\[\frac{(1/x) - (1/y)}{(1/x) + (1/y)}=N\] \[(1/x) - (1/y)=N((1/x) + (1/y))\text{ ; clear fractions}\] \[1 - (x/y)=N(1 + (x/y))\text{ ; clear fractions again}\] \[y - x=N(y + x))\text{ ; divide out the (y+x)}\] \[\frac{y - x}{y+x}=N\text{ ; tada!! :)}\]

OpenStudy (akshay_budhkar):

where did my answer go? did any one delete it?

OpenStudy (amistre64):

i havent seen it ....

OpenStudy (anonymous):

I saw it briefly

OpenStudy (akshay_budhkar):

i replied first when kait said explain please before that

OpenStudy (akshay_budhkar):

then some1 replied so i didnt explain

OpenStudy (akshay_budhkar):

bur my answer is removed

OpenStudy (anonymous):

Thank you guys!!(: & no, i didnt delete it, thats what I was looking at, I couldnt find it. xD

OpenStudy (anonymous):

I often delete answers that have no explanation or justification. It doesn't help the user, it can be easily found using wolframalpha. If you want to help people, help them. Don't give them answers.

OpenStudy (akshay_budhkar):

yes but you should not do that polpak.. i generally answer and ask if the user wants me to explain and then i explain..... or should i change my technique then?

OpenStudy (anonymous):

You should assume that they want an explanation. If they don't they should be using wolframalpha to check their answers (or cheat).

OpenStudy (akshay_budhkar):

okay...

OpenStudy (anonymous):

This is upsetting. But you're fine guys, you're helping me out alot. You're the only three or four people that ever answer my questions anyways.(:

OpenStudy (akshay_budhkar):

kait nothing to be upset of.. All of us are here to help.. All the people ill help you.. I just asked polpak to confirm if he deleted my answer n nothing

OpenStudy (anonymous):

I don't think anyone's upset. Or at least I'm not. Just a bit of administrivia that happened to occur in your question. Not a big deal.

OpenStudy (akshay_budhkar):

right:)